what is
1.-4(x+10)-6=-3(x-2)
2.2/3(x-4)=2x
3.slove for h: A= l times W times H

Answer by Boreal(15235)

1. -4x-40-6=-3x+6
-x=52
x=-52
2. (2/3)x-(8/3)=2x
-(8/3)=(4/3)x
x=-2
3. A=Wh
h=A/W